# Diffusion-Free Dynamics in Rotating Spherical Shell Convection Driven By Internal Heating and Cooling

### v1

@@ Abstract
The bulk properties of convection in stellar and giant planet interiors are often assumed to be independent of the molecular diffusivities, which are very small. By contrast, simulations of this process in rotating, spherical shells, which are typically driven by conductive boundary heat fluxes, generally yield results that depend on the diffusivity. This makes it challenging to extrapolate these simulation results to real objects. However, laboratory models and Cartesian-box simulations suggest diffusion-free dynamics can be obtained if convection is driven using prescribed internal heating and cooling instead of boundary fluxes. Here, we apply this methodology to simulations of Boussinesq, hydrodynamic rotating spherical shell convection. We find that this set-up unambiguously yields diffusion-free behaviour for bulk ‘thermal’ properties of the convection, such as the radial temperature contrast and the convective heat transport. Moreover, the transition from prograde to retrograde equatorial zonal flow is diffusion-free and only depends on the convective Rossby number. The diffusivity dependence of other bulk ‘kinematic’ properties is regime-dependent. In simulations that are rotationally constrained, the convective velocities, and the strength and structure of the zonal flow, are diffusion-dependent, although the zonal flow appears to approach a diffusion-free state for sufficiently high supercriticality. In simulations that are uninfluenced by rotation, or are only influenced by rotation at large scales, diffusion-free convective velocities and zonal flow amplitudes are obtained. The result that many aspects of our idealised simulations are diffusion-free has promising implications for the development of realistic stellar and giant planet convection models that can access diffusion-free regimes.
@@
software: Dedalus v3 (K. J. Burns et al., 2020).
@@ 1 Introduction
Convection in stellar and giant planet interiors is extremely turbulent, which poses a practical barrier to numerical simulation. To retain computational feasibility, simulations of this process are typically conducted in a parameter regime far from reality (A. S. Brun, 2020; P. J. Käpylä et al., 2023), with the results subsequently extrapolated to real objects through scaling analysis and by appealing to theory (J. M. Aurnou et al., 2020). In order to achieve this, it is desirable that the simulations exhibit the same scaling behaviour as would a real planet or star. It is typically assumed that the bulk properties of the convection should not depend on the molecular diffusivities, which are small in real astrophysical environments (K. Julien et al., 2012; A. J. Barker et al., 2014). This assumption underpins the derivation of mixing length theory (MLT; E. Böhm-Vitense, 1958), which has been applied with noteworthy success as a parametrisation of convective heat transport in stellar evolution models (e.g., I. Baraffe et al., 2015).
@@
However, diffusivity independent (i.e., ‘diffusion-free’) dynamics are rarely obtained in numerical simulations. For example, the vast majority of shell convection simulations yield scaling behaviour for the radial heat transport that depends upon the diffusivity (U. R. Christensen, 2002; T. Gastine & J. Wicht, 2012; T. Gastine et al., 2015, 2016; N. A. Featherstone & B. W. Hindman, 2016; B. O’Mara et al., 2016), although a notable exception is presented by T. Gastine et al. (2016), where a subset of their simulations do obtain a diffusion-free scaling for cases with very rapid rotation (paired with moderate supercriticality). The influence of diffusivity in simulations of shell convection also affects their bulk kinematic properties. For example, simulations of stellar convection with high diffusivity have been shown to generate prograde, ‘Solar-like’ equatorial zonal flows (or ‘differential rotation’), but they have a tendency to switch to a regime of retrograde, ‘anti-Solar’ flow as the diffusivities are reduced (T. Gastine et al., 2013, 2014; P. J. Käpylä et al., 2014; N. A. Featherstone & M. S. Miesch, 2015; B. W. Hindman et al., 2020). Moreover, T. Gastine et al. (2016) obtain a diffusivity dependent scaling for the Reynolds number even for simulations that are highly supercritical.
@@
Driving convection using a conductive heat flux through the inner and outer radial boundaries is one cause of diffusivity dependence in numerical simulations. When convection is boundary-driven, thin thermal boundary layers form, and the heat transport through the domain is throttled by the conductive flux through the boundary layer (T. Gastine et al., 2016; see also the review by G. Ahlers et al., 2009). Regarding the differential rotation, the transition from prograde to retrograde motion is understood to occur when the convective Rossby number, which for a fixed rotation rate generally increases as the diffusivities are reduced (in simulations with diffusion-dependent dynamics), exceeds unity (P. A. Gilman, 1977; T. Gastine et al., 2014). M. E. Camisassa & N. A. Featherstone (2022) provide a geometric interpretation for this transition, while other authors (e.g., B. W. Hindman et al., 2020) argue it is directly related to the thinning of the thermal boundary layer. It is interesting to note that simulations of deep stellar convection driven entirely by internal heating and cooling also exhibit diffusivity dependent differential rotation (A. S. Brun & J. Toomre, 2002; M. S. Miesch et al., 2008; M. K. Browning, 2008); however, in these models conductive heat transport is likely still important near the upper boundary, owing to the thinness of the cooling layer.
@@
It can be argued ‘boundary-driven’ simulations are not particularly relevant to real astrophysical environments, as the boundary layers that dominate the dynamics over large swathes of parameter space likely bear little resemblance to those of the real objects (A. J. Barker et al., 2014; B. W. Hindman et al., 2020). This has led some authors to pursue an alternative approach, where convection is driven by radiative heating and cooling (with a specified spatial structure) applied directly to the fluid. In the context of rapidly rotating $f$-plane convection in a Cartesian geometry, A. J. Barker et al. (2014) demonstrate that this approach yields diffusion-free bulk scaling behaviour for both thermal (e.g., the convective heat transport) and kinematic (e.g., the fluctuating kinetic energy) properties. L. K. Currie et al. (2020) generalise this result to the case where the directions of the acceleration due to gravity and the rotational axis are misaligned (as is the case for the non-polar latitudes of a sphere). However, T. Joshi-Hartley et al. (2025) illustrate that it is possible to simultaneously obtain diffusion-free heat transport, but diffusion-dependent convective velocities (using simulations of internally heated and cooled, rotating convection in an idealised Cartesian geometry with $2.5$ dimensions). Additionally, they show that the realisation of diffusion-free dynamics in their model is sensitive to the kinematic boundary condition (free-slip configurations more readily yield diffusion-free behaviour). S. Kazemi et al. (2022) show that diffusion-free heat transport can be obtained in Cartesian simulations of non-rotating convection driven by internal heating and cooling (i.e., a realisation of the ‘ultimate regime’). Finally, laboratory experiments driven by internal heating and cooling have obtained diffusion-free heat transport for both the rotating case and the non-rotating case (S. Lepot et al., 2018; V. Bouillaut et al., 2021; G. Hadjerci et al., 2024).
@@ 2 Methods
Figure 1: Summary of simulations presented in this study. Panel a) shows the input parameters $Ra_{\text{F}}$ and TaTa used for each experiment. Panel b) shows the combination $RaTa^{-\frac{2}{3}}$ as a proxy for the supercriticality (computed using the ‘output’ Rayleigh number given by Equation 7). Panel c) shows ReRe (defined by Equation 8), computed using the r.m.s. fluctuating velocity, as a measure of the degree of turbulence obtained in each simulation. Finally, panel d) shows the output Rossby numbers $Ro_{\text{bulk}}$ and $Ro_{\omega}$ (defined in Equation 9) as a measure of the degree of rotational constraint at large- and small-scales, respectively, plotted against $Ro_{\text{cv,F}}$. The marker colour denotes the dynamical regime occupied by a simulation, and the marker shape denotes the configuration of the differential rotation. In panels a)–c), simulations that share the same $Ro_{\text{cv,F}}$ are connected by lines. Three lines are emphasised in bold, corresponding to $Ro_{\text{cv,F}}=3.11\times 10^{-2},6.70\times 10^{-2},\ \text{and}\ 3.11\times 10^{-1}$ (from bottom to top in each panel). In panel d), the grey dashed lines are proportional to $Ro_{\text{cv,F}}$ and included as an eye guide.
We study the dynamics of Boussinesq, rotating spherical shell convection driven by internal heating and cooling. Our simulations are conducted using the pseudo-spectral code Dedalus (K. J. Burns et al., 2020). All variables are expended in terms of spherical harmonics in the horizontal directions and Chebyshev polynomials in the radial direction using a standard 3/2-dealisaing.
@@ 2.1 Model
where $\boldsymbol{u}=(u_{r},u_{\theta},u_{\phi})$ is the fluid velocity in spherical polar coordinates, $p$ is a pressure, and $T$ is a scaled temperature variable, related to the real temperature by $T=g\alpha\delta T_{\text{real}}$ (where $g$ is the magnitude of gravitational acceleration and $\alpha$ is the coefficient of thermal expansion, and $\delta T_{\text{real}}$ is the Boussinesq temperature perturbation). The symbol $\nu$ denotes the kinematic viscosity and $\kappa$ denotes the thermal conductivity. The dimensionless gravity profile is given by $\tilde{g}(r)=\left(r_{\text{o}}/r\right)^{2}$ where $r_{\text{o}}$ is the outer radius of the convection zone (see below for definition). The unit vector in the radial direction is given by $\mathbf{e}_{r}$, and the rotation vector $\mathbf{\Omega}=\Omega_{0}(\cos\theta,-\sin\theta,\ 0)$, where $\theta$ is co-latitude and $\Omega_{0}$ is the rotation rate.
Convection is driven by a prescribed heating and cooling function $q(r)$, configured so that heat is deposited and removed from regions of depth $\delta$ at the base and top of the simulated domain, respectively. We use the notation $r_{\text{i}}$ and $r_{\text{o}}$ to denote the radii of the inner and outer boundaries of the convection zone (CZ), within which there is no imposed heating or cooling. Mathematically, $q$ is defined (following A. J. Barker et al., 2014; L. K. Currie et al., 2020):
@@ 2.2 Non-dimensionalisation
In all simulations, we set $Pr=1$, $\eta=0.8$, and $\tilde{\delta}=0.2$. From the definitions of $Ra_{\text{F}}$, $Ta$, and $Pr$, it is possible to define a flux-based convective Rossby number (e.g., U. R. Christensen, 2002; U. R. Christensen & J. Aubert, 2006; B. W. Hindman et al., 2020; P. J. Käpylä, 2024), given by:
@@
which, notably, is independent of the diffusivities $\nu$ and $\kappa$. Therefore, for a fixed flux, variation of $Ro_{\text{cv,F}}$ corresponds to varying $\Omega_{0}$, while varying $Ra_{\text{F}}$ and $Ta$ so as to keep $Ro_{\text{cv,F}}$ fixed is equivalent to varying $\nu$ and $\kappa$ (recall $\kappa=\nu$ when $Pr=1$). We note that this definition of $Ro_{\text{cv,F}}$ is closely related to estimates of the Rossby number commonly employed in stellar astronomy (e.g., R. W. Noyes et al., 1984), $Ro\,{\sim}\,1/(\Omega_{0}\tau_{\text{c}})$, assuming that the convective overturning time $\tau_{\text{c}}$ is given by MLT.
@@ 2.3 Summary of experiments and numerical details
Figure 1a shows the input parameters $Ra_{\text{F}}$ and $Ta$, Figure 1b shows $RaTa^{-\frac{2}{3}}$ as a proxy for the supercriticality $Ra/Ra_{\text{c}}$ ($Ra_{\text{c}}\,{\sim}\,Ta^{\frac{2}{3}}$ for large $Ta$; S. Chandrasekhar, 1961), and Figure 1c shows the output Reynolds number to illustrate the degree of turbulence in each experiment. Above, $Ra$ is the usual Rayleigh number, which is an output parameter for a fixed-flux set-up. It is defined as
@@
where for $Ro_{\omega}$, $\omega$ is the r.m.s fluctuating vorticity within the CZ. These two output $Ro$ are intended to indicate the degree of rotational constraint at the dominant scale of convection, and at small scales, respectively. Scaling results obtained from our experiments will show that it is useful to categorise them in terms of three regimes: (1) a ‘rotationally-constrained’ regime defined by $Ro_{\text{cv,F}}<0.1$, for which both large and small scales are rotationally constrained (i.e., $Ro_{\omega}\lesssim 1$ and $Ro_{\text{bulk}}\ll 1$); (2) a ‘rotationally-influenced’ regime defined by $0.1\leq Ro_{\text{cv,F}}<1$, where only the large-scales are rotationally constrained ($Ro_{\omega}\gtrsim 1$ and $Ro_{\text{bulk}}<1$); and (3) a ‘rotationally-uninfluenced’ regime defined by $Ro_{\text{cv,F}}\geq 1$, for which the rotational constraint is lost entirely ($Ro_{\omega}\gg 1$ and $Ro_{\text{bulk}}>1$).

### v2

@@ Abstract
The bulk properties of convection in stellar and giant planet interiors are often assumed to be independent of the molecular diffusivities, which are very small. By contrast, simulations of this process in rotating, spherical shells, which are typically driven by conductive boundary heat fluxes, generally yield results that depend on the diffusivity. This makes it challenging to extrapolate these simulation results to real objects. However, laboratory models and Cartesian-box simulations suggest that diffusion-free dynamics are more readily obtained if convection is driven using prescribed internal heating and cooling instead of boundary fluxes. Here, we apply this methodology to simulations of Boussinesq, hydrodynamic rotating spherical shell convection. We find that this set-up unambiguously yields diffusion-free behaviour for some bulk properties of the convection, such as the radial temperature contrast and the convective heat transport. Moreover, the transition from prograde to retrograde equatorial zonal flow is diffusion-free and only depends on the convective Rossby number. The diffusivity dependence of other bulk properties is regime-dependent. In simulations that are rotationally constrained, the convective velocities, and the strength and structure of the zonal flow, are diffusion-dependent, although the zonal flow appears to approach a diffusion-free state for sufficiently high supercriticality. In simulations that are uninfluenced by rotation, or are only influenced by rotation at large scales, diffusion-free convective velocities and zonal flows are obtained. The result that many aspects of our idealised simulations are diffusion-free has promising implications for the development of realistic stellar and giant planet convection models that can access diffusion-free regimes.
@@
software: Dedalus v3 (2020PhRvR...2b3068B).
@@ 1 Introduction
Convection in stellar and giant planet interiors is extremely turbulent, which poses a practical barrier to numerical simulation. To retain computational feasibility, simulations of this process are typically conducted in a parameter regime far from reality (2020ASSP...57...75B; 2023SSRv..219...58K), with the results subsequently extrapolated to real objects through scaling analysis and by appealing to theory (2020PhRvR...2d3115A). In order to achieve this, it is desirable that the simulations exhibit the same scaling behaviour as would a real planet or star. It is typically assumed that the bulk properties of the convection should not depend on the molecular diffusivities, which are small in real astrophysical environments (2012PhRvL.109y4503J; 2014ApJ...791...13B). This assumption underpins the derivation of mixing length theory (MLT; 1958ZA.....46..108B), which has been applied with noteworthy success as a parametrisation of convective heat transport in stellar evolution models (e.g., 2015A&A...577A..42B).
@@
However, diffusivity independent (i.e., ‘diffusion-free’) dynamics are rarely obtained in numerical simulations. For example, the vast majority of shell convection simulations yield scaling behaviour for the radial heat transport that depends upon the diffusivity (2002JFM...470..115C; 2012Icar..219..428G; 2015JFM...778..721G; 2016JFM...808..690G; 2016ApJ...818...32F; 2016AdSpR..58.1475O), although a notable exception is presented by 2016JFM...808..690G, where a subset of their simulations do obtain a diffusion-free scaling for cases with very rapid rotation (paired with moderate supercriticality). The influence of diffusivity in simulations of shell convection also affects their bulk kinematic properties. For example, simulations of stellar convection with high diffusivity have been shown to generate prograde, ‘Solar-like’ equatorial zonal flows (or ‘differential rotation’), but they have a tendency to switch to a regime of retrograde, ‘anti-Solar’ flow as the diffusivities are reduced (2013Icar..225..156G; 2014MNRAS.438L..76G; 2014A&A...570A..43K; 2015ApJ...804...67F; 2020ApJ...898..120H). Moreover, 2016JFM...808..690G obtain a diffusivity-dependent scaling for the Reynolds number even for simulations that are highly supercritical.
@@
Driving convection using a conductive heat flux through the inner and outer radial boundaries is one cause of diffusivity dependence in numerical simulations. When convection is boundary-driven, thin thermal boundary layers form, and the heat transport through the domain is throttled by the conductive flux through the boundary layer (2016JFM...808..690G; see also the review by 2009RvMP...81..503A). Regarding the differential rotation, the transition from prograde to retrograde motion is understood to occur when the convective Rossby number, which for a fixed rotation rate generally increases as the diffusivities are reduced (in simulations with diffusion-dependent dynamics), exceeds unity (1977GApFD...8...93G; 2014MNRAS.438L..76G). 2022ApJ...938...65C provide a geometric interpretation for this transition, while other authors (e.g., 2020ApJ...898..120H) argue it is directly related to the thinning of the thermal boundary layer. It is interesting to note that simulations of deep stellar convection driven entirely by internal heating and cooling also exhibit diffusivity-dependent differential rotation (2002ApJ...570..865B; 2008ApJ...673..557M; 2008ApJ...676.1262B); however, in these models conductive heat transport is likely still important near the upper boundary, owing to the thinness of the cooling layer.
@@
It can be argued ‘boundary-driven’ simulations are not particularly relevant to real astrophysical environments, as the boundary layers that dominate the dynamics over large swathes of parameter space likely bear little resemblance to those of the real objects (2014ApJ...791...13B; 2020ApJ...898..120H). This has led some authors to pursue an alternative approach, where convection is driven by radiative heating and cooling (with a specified spatial structure) applied directly to the fluid. In the context of rapidly rotating $f$-plane convection in a Cartesian geometry, 2014ApJ...791...13B demonstrate that this approach yields diffusion-free bulk scaling behaviour for both thermal (e.g., the convective heat transport) and kinematic (e.g., the fluctuating kinetic energy) properties. 2020MNRAS.493.5233C generalise this result to the case where the directions of the acceleration due to gravity and the rotational axis are misaligned (as is the case for the non-polar latitudes of a sphere). However, 2025MNRAS.541.2291J illustrate that it is possible to simultaneously obtain diffusion-free heat transport, but diffusion-dependent convective velocities (using simulations of internally heated and cooled, rotating convection in an idealised Cartesian geometry with $2.5$ dimensions). Additionally, they show that the realisation of diffusion-free dynamics in their model is sensitive to the kinematic boundary condition (free-slip configurations more readily yield diffusion-free behaviour). 2022PhRvL.129b4501K show that diffusion-free heat transport can be obtained in Cartesian simulations of non-rotating convection driven by internal heating and cooling (i.e., a realisation of the ‘ultimate regime’). Finally, laboratory experiments driven by internal heating and cooling have obtained diffusion-free heat transport for both the rotating case and the non-rotating case (2018PNAS..115.8937L; 2021PNAS..11805015B; 2024JFM...998A...9H).
@@ 2 Methods
We study the dynamics of Boussinesq, rotating spherical shell convection driven by internal heating and cooling. Our simulations are conducted using the pseudo-spectral code Dedalus (2020PhRvR...2b3068B). All variables are expanded in terms of spherical harmonics in the horizontal directions and Chebyshev polynomials in the radial direction using a standard 3/2 dealiasing.
@@ 2.1 Model
where $\boldsymbol{u}=(u_{r},u_{\theta},u_{\phi})$ is the fluid velocity in spherical polar coordinates, $p$ is a pressure, and $T$ is a scaled temperature variable, related to the real temperature by $T=g\alpha\delta T_{\text{real}}$ (where $g$ is the magnitude of gravitational acceleration and $\alpha$ is the coefficient of thermal expansion, and $\delta T_{\text{real}}$ is the Boussinesq temperature perturbation). The symbol $\nu$ denotes the kinematic viscosity and $\kappa$ denotes the thermal diffusivity. The dimensionless gravity profile is given by $\tilde{g}(r)=\left(r_{\text{o}}/r\right)^{2}$ where $r_{\text{o}}$ is the outer radius of the convection zone (see below for definition). The unit vector in the radial direction is given by $\mathbf{e}_{r}$, and the rotation vector $\mathbf{\Omega}=\Omega_{0}(\cos\theta,-\sin\theta,\ 0)$, where $\theta$ is co-latitude and $\Omega_{0}$ is the rotation rate.
Convection is driven by a prescribed heating and cooling function $q(r)$, configured so that heat is deposited and removed from regions of depth $\delta$ at the base and top of the simulated domain, respectively. We use the notation $r_{\text{i}}$ and $r_{\text{o}}$ to denote the radii of the inner and outer boundaries of the convection zone (CZ), within which there is no imposed heating or cooling. Mathematically, $q$ is defined (following 2014ApJ...791...13B; 2020MNRAS.493.5233C):
@@ 2.2 Non-dimensionalisation
In all simulations, we set $Pr=1$, $\eta=0.8$, and $\tilde{\delta}=0.2$. From the definitions of $Ra_{\text{F}}$, $Ta$, and $Pr$, it is possible to define a flux-based convective Rossby number (e.g., 2002JFM...470..115C; 2006GeoJI.166...97C; 2020ApJ...898..120H; 2024A&A...683A.221K), given by:
@@
which, notably, is independent of the diffusivities $\nu$ and $\kappa$. Therefore, for a fixed flux, variation of $Ro_{\text{cv,F}}$ corresponds to varying $\Omega_{0}$, while varying $Ra_{\text{F}}$ and $Ta$ so as to keep $Ro_{\text{cv,F}}$ fixed is equivalent to varying $\nu$ and $\kappa$ (recall $\kappa=\nu$ when $Pr=1$). We note that this definition of $Ro_{\text{cv,F}}$ is closely related to estimates of the Rossby number commonly employed in stellar astronomy (e.g., 1984ApJ...279..763N), $Ro\,{\sim}\,1/(\Omega_{0}\tau_{\text{c}})$, assuming that the convective overturning time $\tau_{\text{c}}$ is given by MLT.
@@ 2.3 Summary of experiments and numerical details
Figure 1: Summary of simulations presented in this study. Panel a) shows the input parameters $Ra_{\text{F}}$ and TaTa used for each experiment. Panel b) shows the combination $RaTa^{-\frac{2}{3}}$ as a proxy for the supercriticality (computed using the ‘output’ Rayleigh number given by Equation 7). Panel c) shows ReRe (defined by Equation 8), computed using the r.m.s. fluctuating velocity, as a measure of the degree of turbulence obtained in each simulation. Finally, panel d) shows the output Rossby numbers $Ro_{\text{bulk}}$ and $Ro_{\omega}$ (defined in Equation 9) as a measure of the degree of rotational constraint at the largest spatial scale, and smaller scales, respectively, plotted against $Ro_{\text{cv,F}}$. The marker colour denotes the dynamical regime occupied by a simulation, and the marker shape denotes the configuration of the differential rotation. In panels a)–c), simulations that share the same $Ro_{\text{cv,F}}$ are connected by lines. Three lines are emphasised in bold, corresponding to $Ro_{\text{cv,F}}=3.11\times 10^{-2},6.70\times 10^{-2},\ \text{and}\ 3.11\times 10^{-1}$ (from bottom to top in each panel). In panel d), the grey dashed lines are proportional to $Ro_{\text{cv,F}}$ and included as an eye guide.
@@
Figure 1a shows the input parameters $Ra_{\text{F}}$ and $Ta$, Figure 1b shows $RaTa^{-\frac{2}{3}}$ as a proxy for the supercriticality $Ra/Ra_{\text{c}}$ ($Ra_{\text{c}}\,{\sim}\,Ta^{\frac{2}{3}}$ for large $Ta$; 1961hhs..book.....C), and Figure 1c shows the output Reynolds number to illustrate the degree of turbulence in each experiment. Above, $Ra$ is the usual Rayleigh number, which is an output parameter for a fixed-flux set-up. It is defined as
@@
where for $Ro_{\omega}$, $\omega$ is the r.m.s fluctuating vorticity within the CZ. These two output $Ro$ are intended to indicate the degree of rotational constraint at the largest spatial scales, and at small scales, respectively (e.g., 2021PNAS..11822518V). We note that $Ro_{\omega}$ returns very similar values to the ‘local Rossby number’ $Ro_{l}=Ro_{\text{bulk}}\frac{\ell}{\pi}$, where $\ell$ is the mean spherical harmonic degree obtained from the (mid-plane) kinetic energy spectrum (2014MNRAS.438L..76G; not shown).
Scaling results obtained from our experiments will show that it is useful to categorise them in terms of three regimes: (1) a ‘rotationally-constrained’ regime defined by $Ro_{\text{cv,F}}<0.1$, for which all spatial scales are rotationally constrained (i.e., $Ro_{\omega}\lesssim 1$ and $Ro_{\text{bulk}}\ll 1$); (2) a ‘rotationally-influenced’ regime defined by $0.1\leq Ro_{\text{cv,F}}<1$, where the largest scales are rotationally constrained, but (at least some) smaller scales are not ($Ro_{\omega}\gtrsim 1$ and $Ro_{\text{bulk}}<1$); and (3) a ‘rotationally-uninfluenced’ regime defined by $Ro_{\text{cv,F}}\geq 1$, for which the rotational constraint is lost entirely ($Ro_{\omega}\gg 1$ and $Ro_{\text{bulk}}>1$). From Figure 1d, it is clear that $Ro_{\omega}$ is not diffusion-free (i.e., there is a spread in $Ro_{\omega}$ for a given $Ro_{\text{cv,F}}$). This is because $Ro_{\omega}$ is sensitive to the dissipation scale, which will necessarily be influenced by the diffusivity. However, the dependence of $Ro_{\omega}$ on the diffusivity should not be taken to indicate that the transition between the rotationally-constrained and rotationally-influenced regimes is necessarily diffusivity-dependent; our results will show that these regimes are most accurately demarcated by $Ro_{\text{cv,F}}$ (which is diffusion-free). Instead, it suggests that $Ro_{\omega}$ (and $Ro_{l}$) is an imperfect measure of the influence of rotation on smaller scales within the inertial range, which, in general, may or may not be sensitive to the diffusivity for a given, intermediate $\ell$.